A father is 23 years older than his son. The sum of their ages is now 69. How old is their son?


Found 2 solutions by addingup, Alan3354:
Answer by addingup(3677)

f = s+23
f+s = 69
s+23+s = 69
2s+23 = 69
2s = 46
s = 23 the son is 23 and the father 69-23 = 46

Answer by Alan3354(69443)

A father is 23 years older than his son. The sum of their ages is now 69. How old is their son?
===========
F - S = 23
F + S = 69
------------------ Add